Find the number of milliliters of a solution which contains 0.2000 mole BaCl2 in 500 ml solution that is needed to precipitate SO4-2 as BaSO4 from 3 mM Al2(SO4)3?
1
Expert's answer
2020-05-13T14:19:23-0400

Al2(SO4)3 + 3BaCl2 = 3BaSO4 + 2AlCl3

C(BaCl2) = 0.2/0.5 = 0.4 M

n(SO42-) = 3*3/1000 = 0.009 mol

n(BaCl2) = 0.009 mol

V(BaCl2) = 0.009/0.4 = 0.0225 L = 22.5 mL